# Fast offline decoding with local message-passing automata

### v1

@@ 4.2.4 Erosion of multiple clusters
Consider a state initialized with trivial messages $\mathbf{m}={\mathbf{0}}$. For $v>2$, there exists choices of parameters $w_{0},b_{0},n$ such that for all $k\leq k_{L}$, worldlines of anyons that originate in a $k$-cluster $\mathcal{C}_{k}$ centered at position $\mathbf{r}_{\mathcal{C}_{k}}$ are contained in a spacetime rectangle ~$\widetilde{\mathcal{R}}_{\mathcal{C}_{k}}$ based at $(\mathbf{r}_{\mathcal{C}_{k}},0)$ whose height $T_{k}$ and width $W_{k}$ are upper bounded as
@@
~$\widetilde{\mathcal{R}}_{\mathcal{C}_{k}}\cap\widetilde{\mathcal{R}}_{\mathcal{C}^{\prime}_{k}}=\varnothing.$ (51)
@@
where as before $d_{v}=\frac{3}{2}(1+1/v)$. The erosion analysis guarantees that ~$\widetilde{\mathcal{R}}_{\mathcal{C}_{0}}\cap\widetilde{\mathcal{R}}_{\mathcal{C}_{0}^{\prime}}=\varnothing$, and that anyons originating from $\mathcal{C}_{0}$ never annihilate with anyons from any other 0-cluster.
Figure 7: An illustration of the causal relationships between different regions of spacetime appearing during the proof of Lemma 4.7 in one dimension. Blue lines indicate signals. Anyons in $k$-clusters would be confined to the spacetime boxes marked $\mathcal{R}_{\mathcal{C}_{k}}$ in the absence of message-induced interactions between clusters; the Lemma shows that in the presence of interactions they are instead confined to larger spacetime rectangles ~$\widetilde{\mathcal{R}}_{\mathcal{C}_{k}}$ of height $T_{k}$, with ~$\widetilde{\mathcal{R}}_{\mathcal{C}_{k}}\cap\widetilde{\mathcal{R}}_{\mathcal{C}_{k}^{\prime}}=\varnothing$ for distinct $k$-clusters $\mathcal{C}_{k},\mathcal{C}_{k^{\prime}}$. $t_{\sf rec}$ is an upper bound on the latest time at which the rightmost anyon in a $k$-cluster $\mathcal{C}_{k}$ can receive right-moving signals from another anyon in $\mathcal{C}_{k}$, and $t_{\sf rev}$ is an upper bound on the latest time at which the rightmost anyon initially in $\mathcal{C}_{k}$ can move right. The parameters $w_{0},b_{0},n$ are chosen such that at time $t_{\sf rev}$, the rightmost anyon originally in $\mathcal{C}_{k}$ has to begin moving left before receiving signals from a distinct $k$ cluster $\mathcal{C}^{\prime}_{k}$; this necessitates that the orange dot in the figure lie to the left of the green dot. In higher dimensions, the linear dimensions of the spacetime rectangles $\mathcal{R}_{\mathcal{C}_{k}},\widetilde{\mathcal{R}}_{\mathcal{C}_{k}}$ are larger due to the extra $\mathscr{P}^{x}$ regions occupied by anyons during cluster erosion, c.f. Corollary 4.5.
We now assume that for all $k^{\prime}<k$, $T_{k^{\prime}}\leq Aw_{k^{\prime}},W_{k^{\prime}}\leq Bw_{k^{\prime}}$, that ~$\widetilde{\mathcal{R}}_{\mathcal{C}_{k^{\prime}}}\cap\widetilde{\mathcal{R}}_{\mathcal{C}_{k^{\prime}}^{\prime}}=\varnothing$ for distinct $k^{\prime}$-clusters $\mathcal{C}_{k^{\prime}},\mathcal{C}^{\prime}_{k^{\prime}}$, and that anyons in $\mathcal{C}_{k^{\prime}}$ only annihilate with other anyons in $\mathcal{C}_{k^{\prime}}$. Our aim is to show that there exist $O(1)$ choices of $A,B$ such that the induction step holds, provided $w_{0},b_{0},n$ are chosen appropriately.
@@
In the following we will fix coordinates where $\mathbf{r}_{\mathcal{C}_{k}}={\mathbf{0}}$. Define $t_{\sf rec}$ and $d_{\sf in}(t)$ as in the proof of Corollary 4.6, with $t_{\sf rec}<w_{k}/(v-1)$ [^20] It is $<$ here and not $\leq$ since anyons in $\mathcal{C}_{k}$ cannot actually begin to move into $\Lambda\setminus B_{{\mathbf{0}}}^{(\infty)}(W/2)$ until a time of at least $b_{0}/v$ (as we initialize the system with trivial messages). and $d_{\sf in}(t)<w_{k}+t_{\sf rec}+2(t-t_{\sf rec})$. Define the regions
@@
Since the closest worldlines in the past lightfronts of anyons in $\mathscr{B}_{\mathcal{C}_{k}}(t)$ belong to anyons which originated in $\mathcal{C}_{k}$ when $d_{\sf out}(t)>d_{\sf in}(t)$, the satisfaction of (55) at $t=t_{\sf rev}$ will guarantee that anyons from $\mathcal{C}_{k}$ will not annihilate with anyons from any distinct $k$-cluster $\mathcal{C}_{k}^{\prime}$, showing that ~$\widetilde{\mathcal{R}}_{\mathcal{C}_{k}}\cap\widetilde{\mathcal{R}}_{\mathcal{C}^{\prime}_{k}}=\varnothing$ as desired. Using the above bound for $t_{\sf rev}$, this happens if
@@ 5.1 Preliminaries
$p_{\sf log}^{(\mathcal{E})}=\frac{1}{2}\max_{\rho_{\sf log}}\lim_{t\rightarrow\infty}||\mathcal{D}^{t}(\mathcal{E}(\rho_{\sf log}))_{\sf qu}-\rho_{\sf log}||_{1}.$ (65)
@@
The discussion in Sec. 2.2 means that we may without loss of generality take ${\rho_{\sf log}}$ to be the $+1$ eigenstate of each logical $Z$ operator, and (as before) take $\mathcal{E}$ to only apply Kraus operators built from tensor products of $X$ operators; these choices will be made in what follows, and we will consequently omit the ${\rho_{\sf log}}$ in $\rho\sim\mathcal{E},{\rho_{\sf log}}$.
@@ 5.2.2 Bounding $T_{\sf dec}^{(\mathcal{E})}$
where the second line on the RHS requires $\lambda>1$ (with $22D\lambda/\log_{n}$ 2 at e.g. $n=13$ setting a bound on the constant $c$ above).
@@ 5.3 General comments on decoding times
below threshold for noise channels $\mathcal{E}$ with local spatial correlations. Our threshold proof gave weaker bounds, but in a more general setting (for $\mathcal{E}$ with non-local correlations). One may naturally wonder how tight the bounds of our threshold result are: could a refinement of our proof be used to improve the guaranteed scaling to (87) for all $p$-bounded $\mathcal{E}$? The following proposition shows that the answer is negative: there are indeed (spatially correlated) $p$-bounded error models for which $p_{\sf log}^{(\mathcal{E})}$ is only a stretched exponential in $L$, and for which $T_{\sf dec}^{(\mathcal{E})}$ scales as $(\log L)^{\beta}$ for some $\beta>1$. This implies that a proof that $T_{\sf dec}^{(\mathcal{E})}$ scales as in (87) for spatially local $\mathcal{E}$ will necessitate techniques that specifically take advantage of the lack of $\mathcal{E}$’s spatial correlations. [^26] The fractal-like error patterns of the following proposition will likely dominate the scaling of $p_{\sf log}^{(\mathcal{E})}$ at small enough $p$—even for local $\mathcal{E}$—but it is still possible for them to be too rare to influence $T_{\sf dec}^{(\mathcal{E})}$.
Proposition 5.3.
@@
with $\rho_{0}=(|0\rangle\langle 0|)^{\otimes L}$. We claim that for any update speed $v\geq 3$, there exist choices of $k<n$ such that the decoder outputs $\rho_{1}=(|1\rangle\langle 1|)^{\otimes L}$ when run for long enough on input state $\rho$ [^27] For simplicity, we use the deterministic feedback rule, where an anyon does not move along ${\mathbf{\hat{a}}}$ if the forces it receives are balanced along the ±$\pm{\mathbf{\hat{a}}}$ directions. To ensure that the logical states are the only steady states of the decoder, this mandates that the system size (and hence $n$) be odd.
For simplicity of presentation, we will specify to $k=n-1$ and momentarily take $v=\infty$. We then prove the that the decoder maps $\rho_{n,k}$ to $(|1\rangle\langle 1|)^{\otimes L}$ after a time linear in $L$ if $n\geq 5$. To ease the notation, we will use $\mathscr{S}$ as shorthand for $\mathscr{S}_{n,n-1}$.
@@
This shows that for $\rho\sim\mathcal{E}_{i.i.d}$, there is a matching of the anyons in $\rho$ whose edge lengths are all less than $b(\log L)^{1/D}$ w.h.p. Whether or not there is a perfect matching with the same bound on the largest edge length is a different question—since the largest separation between two anyons paired by the operators created by the noise goes like $\log(L)$ w.h.p, independent of $D$—although the following footnote argues that there is. [^29] As just mentioned, there always exists a perfect matching in $\rho$ which pairs anyons up according to the way in which they were created by the noise, and the longest edge length in this perfect matching goes as $\log(L)$ w.h.p. Given an anyon pair at coordinates $\mathbf{r}_{1},\mathbf{r}_{2}$ with $||\mathbf{r}_{1}-\mathbf{r}_{2}||\sim\log(L)$, we can however w.h.p find a “bridge” of smaller anyon pairs connecting the two anyons in the large pair, in the sense that there exist anyon pairs at coordinates $\mathbf{r}_{i,1},\mathbf{r}_{i,2}$, $i=1,\dots,n$ with $n\lesssim(\log L)^{1-1/D}$, $||\mathbf{r}_{1}-\mathbf{r}_{1,1}||,||\mathbf{r}_{i,\alpha}-\mathbf{r}_{i+1,\beta}||,||\mathbf{r}_{n,2}-\mathbf{r}_{2}||<b(\log L)^{1/D}$ w.h.p. Flipping the perfect matching along the anyons in this bridge then gives a perfect matching where the edge of length $\sim\log(L)$ has been substituted for one of length at most $\sim(\log L)^{1/D}$. Repeating this process for all $\sim\log(L)$-sized pairs then suggests the existence of a perfect matching with largest edge length $\sim(\log L)^{1/D}$. Nevertheless, constructing a perfect matching with this bound on the largest edge length appears to require nonlocal information, and we conjecture that $\log(L)$ is the best one can do for i.i.d noise with a local decoder. Since any noise model with quickly decaying spatial correlations should roughly speaking renormalize to i.i.d noise at large enough (but still $O(L^{0})$) distances, we conjecture the following:
@@ 6.1 Desynchronization with marching soldiers
$\gamma_{i}=||L_{i}|\psi(t)\rangle||^{2}=\frac{1}{\mu}\prod_{\mathbf{r}^{\prime}\in B_{\mathbf{r}}(\Delta)}\delta_{m_{\mathbf{r}^{\prime}},\mathcal{A}_{\sf desync}(\mathbf{m}^{\prime},{\bm{\sigma}}^{\prime})[m_{\mathbf{r}^{\prime}}]}\langle\psi(t)|\left(|\mathbf{m}^{\prime}{\bm{\sigma}}^{\prime}\rangle\langle\mathbf{m}^{\prime}{\bm{\sigma}}^{\prime}|\otimes\Pi|_{{\bm{\sigma}}|_{B_{\mathbf{r}}(\Delta)}}\right)|\psi(t)\rangle.$ (106)
@@
Our construction for producing a Lindbladian decoder from a synchronous one is based on the “marching soldiers scheme” of Refs. [28, 36, 29], which as we will see can be implemented at the cost of increasing the number of classical control variables stored at each site by a constant factor. The basic idea is to use $\mathcal{D}_{\sf desync}$ to simulate the computation performed by a given synchronous decoder $\mathcal{D}$. This is done by having each site keep track of how many automaton updates it has performed, and to then require that the number of automaton updates performed by neighboring sites cannot differ by more than 1: if a jump operator applied by $\mathcal{D}_{\sf desync}$ attempts to update a site which would violate this rule, it instead acts trivially. This ensures that no site is “left behind” during the evolution, with nearby sites guaranteed to have made progress on $\mathcal{D}$’s computation which differs by at most a single time step (see Fig. 8 for a schematic). We let each site store the values of its control variables for both the current and previous simulation time steps, and $\mathcal{D}_{\sf desync}$ may then draw on this information to update any updateable site in the same way that $\mathcal{D}$ would. Implementing this scheme is slightly complicated by the fact that the feedback operators of $\mathcal{D}$ will generically modify the values of syndromes on multiple sites, before these sites have had a chance to be updated. This necessitates also storing several variables related to the recent history of syndrome measurements, in a manner that will be explained below.
@@
Furthermore, if ${\sf lag}_{\mathbf{r}\rightarrow\mathbf{r}^{\prime}}\leq 0\,\,\forall\,\,\mathbf{r}^{\prime}\in\widetilde{B}_{\mathbf{r}}(2\Delta)$—so that a proposed automaton update at $\mathbf{r}$ will be accepted—$\mathcal{A}_{\sf desynch}$ simultaneously performs the following updates:

### v4

@@ 4.2.4 Erosion of multiple clusters
Consider a state initialized with trivial messages $\mathbf{m}={\mathbf{0}}$. For $v>2$, there exists choices of parameters $w_{0},b_{0},n$ such that for all $k\leq k_{L}$, worldlines of anyons that originate in a $k$-cluster $\mathcal{C}_{k}$ centered at position $\mathbf{r}_{\mathcal{C}_{k}}$ are contained in a spacetime rectangle $\widetilde{\mathcal{R}}_{\mathcal{C}_{k}}$ based at $(\mathbf{r}_{\mathcal{C}_{k}},0)$ whose height $T_{k}$ and width $W_{k}$ are upper bounded as
@@
$\widetilde{\mathcal{R}}_{\mathcal{C}_{k}}\cap\widetilde{\mathcal{R}}_{\mathcal{C}^{\prime}_{k}}=\varnothing.$ (51)
@@
where as before $d_{v}=\frac{3}{2}(1+1/v)$. The erosion analysis guarantees that $\widetilde{\mathcal{R}}_{\mathcal{C}_{0}}\cap\widetilde{\mathcal{R}}_{\mathcal{C}_{0}^{\prime}}=\varnothing$, and that anyons originating from $\mathcal{C}_{0}$ never annihilate with anyons from any other 0-cluster.
Figure 7: An illustration of the causal relationships between different regions of spacetime appearing during the proof of Lemma 4.7 in one dimension. Blue lines indicate signals. Anyons in $k$-clusters would be confined to the spacetime boxes marked $\mathcal{R}_{\mathcal{C}_{k}}$ in the absence of message-induced interactions between clusters; the Lemma shows that in the presence of interactions they are instead confined to larger spacetime rectangles $\widetilde{\mathcal{R}}_{\mathcal{C}_{k}}$ of height $T_{k}$, with $\widetilde{\mathcal{R}}_{\mathcal{C}_{k}}\cap\widetilde{\mathcal{R}}_{\mathcal{C}_{k}^{\prime}}=\varnothing$ for distinct $k$-clusters $\mathcal{C}_{k},\mathcal{C}_{k^{\prime}}$. $t_{\sf rec}$ is an upper bound on the latest time at which the rightmost anyon in a $k$-cluster $\mathcal{C}_{k}$ can receive right-moving signals from another anyon in $\mathcal{C}_{k}$, and $t_{\sf rev}$ is an upper bound on the latest time at which the rightmost anyon initially in $\mathcal{C}_{k}$ can move right. The parameters $w_{0},b_{0},n$ are chosen such that at time $t_{\sf rev}$, the rightmost anyon originally in $\mathcal{C}_{k}$ has to begin moving left before receiving signals from a distinct $k$ cluster $\mathcal{C}^{\prime}_{k}$; this necessitates that the orange dot in the figure lie to the left of the green dot. In higher dimensions, the linear dimensions of the spacetime rectangles $\mathcal{R}_{\mathcal{C}_{k}},\widetilde{\mathcal{R}}_{\mathcal{C}_{k}}$ are larger due to the extra $\mathscr{P}^{x}$ regions occupied by anyons during cluster erosion, c.f. Corollary 4.5.
We now assume that for all $k^{\prime}<k$, $T_{k^{\prime}}\leq Aw_{k^{\prime}},W_{k^{\prime}}\leq Bw_{k^{\prime}}$, that $\widetilde{\mathcal{R}}_{\mathcal{C}_{k^{\prime}}}\cap\widetilde{\mathcal{R}}_{\mathcal{C}_{k^{\prime}}^{\prime}}=\varnothing$ for distinct $k^{\prime}$-clusters $\mathcal{C}_{k^{\prime}},\mathcal{C}^{\prime}_{k^{\prime}}$, and that anyons in $\mathcal{C}_{k^{\prime}}$ only annihilate with other anyons in $\mathcal{C}_{k^{\prime}}$. Our aim is to show that there exist $O(1)$ choices of $A,B$ such that the induction step holds, provided $w_{0},b_{0},n$ are chosen appropriately.
@@
In the following we will fix coordinates where $\mathbf{r}_{\mathcal{C}_{k}}={\mathbf{0}}$. Define $t_{\sf rec}$ and $d_{\sf in}(t)$ as in the proof of Corollary 4.6, with $t_{\sf rec}<w_{k}/(v-1)$ [^19] It is << here and not $\leq$ since anyons in $\mathcal{C}_{k}$ cannot actually begin to move into $\Lambda\setminus B_{{\mathbf{0}}}^{(\infty)}(W/2)$ until a time of at least $b_{0}/v$ (as we initialize the system with trivial messages). and $d_{\sf in}(t)<w_{k}+t_{\sf rec}+2(t-t_{\sf rec})$. Define the regions
@@
Since the closest worldlines in the past lightfronts of anyons in $\mathscr{B}_{\mathcal{C}_{k}}(t)$ belong to anyons which originated in $\mathcal{C}_{k}$ when $d_{\sf out}(t)>d_{\sf in}(t)$, the satisfaction of (55) at $t=t_{\sf rev}$ will guarantee that anyons from $\mathcal{C}_{k}$ will not annihilate with anyons from any distinct $k$-cluster $\mathcal{C}_{k}^{\prime}$, showing that $\widetilde{\mathcal{R}}_{\mathcal{C}_{k}}\cap\widetilde{\mathcal{R}}_{\mathcal{C}^{\prime}_{k}}=\varnothing$ as desired. Using the above bound for $t_{\sf rev}$, this happens if
@@ 5.1 Preliminaries
[…]
@@
The discussion in sec. 2.2 means that we may without loss of generality take $\rho_{\sf log}$} to be the +1+1 eigenstate of each logical $Z$ operator, and (as before) take $\mathcal{E}$ to only apply Kraus operators built from tensor products of $X$ operators; these choices will be made in what follows, and we will consequently omit the $\rho_{\sf log}$} in $\rho\sim\mathcal{E},{\rho_{\sf log}}$.
@@ 5.2.2 Bounding $T_{\sf dec}^{(\mathcal{E})}$
where the second line on the RHS requires $\lambda>1$ (with $2D\lambda/\log_{n}2$ at e.g. $n=13$ setting a bound on the constant $c$ above).
@@ 5.3 Gerrymandering
below threshold for noise channels $\mathcal{E}$ with local spatial correlations. Our threshold proof gave weaker bounds, but in a more general setting (for $\mathcal{E}$ with non-local correlations). One may naturally wonder how tight the bounds of our threshold result are: could a refinement of our proof be used to improve the guaranteed scaling to (87) for all $p$-bounded $\mathcal{E}$?
The following proposition shows that the answer is negative: there are indeed (highly spatially correlated) $p$-bounded error models for which $p_{\sf log}^{(\mathcal{E})}$ is only a stretched exponential in $L$, and for which $T_{\sf dec}^{(\mathcal{E})}$ scales as $(\log L)^{\beta}$ for some $\beta>1$. This is possible because the message-passing decoder produces logical errors on certain noise events whose weight is sublinear in $L$, which will be demonstrated shortly by showing the existence of certain sparse fractal-like patterns of errors that are guaranteed to cause logical failures. We will refer to this phenomenon as Gerrymandering, since the logical failures arise from a dilute collection of strategiclally-placed errors that sway the outcome of the global error correction operation. Gerrymandering is ubiquitous in error correction algorithms based (either explicitly or implicitly) on clustering; see also the discussions in Refs. [51, 52, 41, 15].
Gerrymandering implies that proving a scaling of $T_{\sf dec}^{(\mathcal{E})}$ as in (87) for spatially local $\mathcal{E}$ will necessitate techniques that specifically take advantage of the lack of $\mathcal{E}$’s spatial correlations. [^25] Gerrymandered configurations can dominate the scaling of $p_{\sf log}^{(\mathcal{E})}$ at very small pp—even for local $\mathcal{E}$—but they are usually too entropically disfavored to influence $T_{\sf dec}^{(\mathcal{E})}$.
Proposition 5.3 (Gerrymandering).
@@
with $\rho_{0}=(|0\rangle\langle 0|)^{\otimes L}$. We claim that for any update speed $v\geq 3$, there exist choices of $k<n$ such that the decoder outputs $\rho_{1}=(|1\rangle\langle 1|)^{\otimes L}$ when run for long enough on input state $\rho$. [^26] For simplicity, we use the deterministic feedback rule, where an anyon does not move along $^{\mathbf{\hat{a}}}$ if the forces it receives are balanced along the $\pm{\mathbf{\hat{a}}}$ directions. To ensure that the logical states are the only steady states of the decoder, this mandates that the system size (and hence $n$) be odd.
For simplicity of presentation, we will specify to k=n1k=n-1 and momentarily take $v=\infty$. We then prove the that the decoder maps $\rho_{n,k}$ to $(|1\rangle\langle 1|)^{\otimes L}$ after a time linear in $L$ if $n\geq 5$. To ease the notation, we will use $\mathscr{S}$ as shorthand for $\mathscr{S}_{n,n-1}$.
@@
[…]
@@
5.4 Extension to arbitrary topological codes
In this short subsection, we show that the message-passing decoder has thresholds for all topological codes $\mathcal{C}$, not just the surface code. In particular, acheiving a threshold for non-Abelian anyon theories does not require modifying the message-passing achitecture in any way, regardless of whether or not they are acyclic (although making $\mathcal{C}$-dependent modifications may of course increase the numerical value of threshold).
We first define how our decoder operates for a generic anyon theory $\mathcal{C}$. Following most treatments of non-Abelian decoders in the literature (e.g. [53, 54, 55, 52]), we work in a phenomenological setting in which anyons live on the sites of the lattice, and charge-neutral pairs of anyons are created by the noise according to a $p$-bounded noise distribution on the links of the lattice (see e.g. [56] for a discussion of how a general qubit-based noise model may be converted to a phenomenological anyon-based one). We assume the ability to make topological charge measurements at each site, [^29] Performing measurements by braiding only unambiguously resolves the anyon type in modular anyon theories; for non-modular theories, the anyon type is resolved only up to elements of the Muger center of the associated tensor category. Since the anyons in the center cannot implement logical operations on a torus, we will not care about keeping track of them, and they can be ignored as far as decoding is concerned. and also assume the ability to deterministically move anyons from one lattice site to another through the application of appropriate string operators (which in general require linear-depth circuits to be implimented, but this point is unimportant for us).
With the ability to measure topological charge and move anyons, we may impliment decoding in precisely the same way as was done for the surface code. In particular, nontrivial anyons always act as message sources, and the messages they send are independent of their topological charge. [^30] We chose to use separate messages for charges and fluxes in our discussion of the surface code, but this is not actually necessary, and by the arguments in this paragraph, ignoring the distinction between them can be done without compromising the existence of a threshold (although doing so will certainly result in a smaller value of $p_{c}$). The decoder then moves anyons in the direction of the smallest-valued recieved message, and topological charges are re-measured at each time step after motion has occured. The non-Abelian nature of the anyons involved does not compromise the existence of a threshold because, at small enough $p$, the decoder has a very high probability of fusing all anyons within a given noise cluster together. Since the anyons in a given noise cluster were created from the vacuum (by the noise), they are as a whole topologically neutral, and fusing them together (in any order) is guaranteed to produce the vacuum, independent of any mutual braiding events among them which may have occured along the way (for a similar discussion in the context of RG-based decoders, see [52, 57]). This argument thus implies
Corollary 5.5 (message-passing thresholds for arbitrary topological codes).
The statements of theorem 5.1 hold independent of $\mathcal{C}$.
@@ 6.1 Desynchronization with marching soldiers
[…]
@@
Our construction for producing a Lindbladian decoder from a synchronous one is based on the “marching soldiers scheme” of Refs. [28, 37, 29], which as we will see can be implemented at the cost of increasing the number of classical control variables stored at each site by a constant factor. The basic idea is to use $\mathcal{D}_{\sf desync}$ to simulate the computation performed by a given synchronous decoder $\mathcal{D}$. This is done by having each site keep track of how many automaton updates it has performed, and to then require that the number of automaton updates performed by neighboring sites cannot differ by more than 1: if a jump operator applied by $\mathcal{D}_{\sf desync}$ attempts to update a site which would violate this rule, it instead acts trivially. This update rule is directly analogous to the process by which surface growth occurs in the restricted solid-on-solid model [30], with the constraint ensuring that no site is “left behind” during the evolution: nearby sites are guaranteed to have made progress on $\mathcal{D}$’s computation by an amount that differs by at most a single time step (see fig. 8 for a schematic). We let each site store the values of its control variables for both the current and previous simulation time steps, and $\mathcal{D}_{\sf desync}$ may then draw on this information to update any updateable site in the same way that $\mathcal{D}$ would. Implementing this scheme is slightly complicated by the fact that the feedback operators of $\mathcal{D}$ will generically modify the values of syndromes on multiple sites, before these sites have had a chance to be updated. This necessitates also storing several variables related to the recent history of syndrome measurements, in a manner that will be explained below.
@@
[…]